Problem 1 (6 pointe study of the amount 10 wee The manager of Fred s Grocery is making a money customers spend in the store. A sample of kda the tollowing amounta epens and 15 Saturday morning shoppers revealed Meekday $18.88 turday $11.54 24-76 Average of Weekdays: 50.347 Average of Saturday: 57.36 Standará Deviation of Reekday: 24.258 Standard Deviation of Saturday: 22.344 24.33 27.26 35.79 42.31 53.77 62.94 73.59 76.51 88.09 5.78 36.87 46.78 56.04 58.45 60.98 62.67 66.89 71-65 75.61 81.87 84-71 85.80 than weekday shoppera. Use the Student t Test and a 5 significance level. Owner Fred Snead contends that Saturday morning shoppers spend more a. State the null and alternate hypotheses HI : Ho: b. State the decision rule. c. Compute the value of the test statistic
Answer #1

Solution:

Let mu_{1} be the mean for Saturday

Let μ2 be the mean for Weekday

a. State the null and alternative hypotheses.

Answer:

oldsymbol{H_{0}:mu_{1}=mu_{2}}

oldsymbol{H_{1}:mu_{1}>mu_{2}}

b. State the decision rule.

Answer: To find the decision rule, we need to first find the right-tailed critical value at 0.05 significance level and for df=n1 + n2-2=10+15-2-23 . Using the t distribution table, we have:

tcriticalー1.714

Therefore, the decision rule is:

Reject the null hypothesis if small oldsymbol{t>1.714}

c. Compute the value of the test statistic.

Answer:

C1-2

ni + n2 - 2

(15- 1)22.3442 (10 - 1)24.2582 1510- 2

  large =23.1118

7.36 50.347 23. 1 1 18yT+το

large =rac{7.013}{9.4354}

large =0.743

Therefore, the test statistic is t-0.743